OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (5,-5); 6x-7y=8 The equation of the line is y=?

OpenStudy (anonymous):

6x-7y = 8 7y = 6x-8 y = (6/7) x - 8/7 slope of this line is 6/7 now find equation of your line: y-y1 = (6/7) (x - x1) y - -5 = (6/7)(x - 5) y + 5 = (6/7)x - 30/7 y = (6/7)x - 65/7

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

yw